# Boosting Omnidirectional Stereo Matching with a Pre-trained Depth Foundation Model

### v1

@@ Abstract
Omnidirectional depth perception is essential for mobile robotics applications that require scene understanding across a full 360° field of view. Camera-based setups offer a cost-effective option by using stereo depth estimation to generate dense, high-resolution depth maps without relying on expensive active sensing. However, existing omnidirectional stereo matching approaches achieve only limited depth accuracy across diverse environments, depth ranges, and lighting conditions, due to the scarcity of real-world data. We present DFI-OmniStereo, a novel omnidirectional stereo matching method that leverages a large-scale pre-trained foundation model for relative monocular depth estimation within an iterative optimization-based stereo matching architecture. We introduce a dedicated two-stage training strategy to utilize the relative monocular depth features for our omnidirectional stereo matching before scale-invariant fine-tuning. DFI-OmniStereo achieves state-of-the-art results on the real-world Helvipad dataset, reducing disparity MAE by approximately 16% compared to the previous best omnidirectional stereo method. More details are available at https://vita-epfl.github.io/DFI-OmniStereo-website/.
@@ I INTRODUCTION
Figure 1: Overview of our proposed omnidirectional stereo matching approach DFI-OmniStereo. Given a pair of equirectangular images captured by two vertically stacked omnidirectional cameras, our method integrates a large-scale pre-trained monocular relative depth foundation model into an iterative stereo matching approach. DFI-OmniStereo improves disparity estimation accuracy, significantly outperforming the previous state-of-the-art method.
@@
Specifically, our contributions are as follows:
  1. 1.
    We leverage a large-scale pre-trained monocular depth foundation model as a feature extractor integrated into an iterative optimization-based stereo matching architecture;
  2. 2.
    we introduce a two-stage training strategy to adapt monocular foundation model features to the omnidirectional stereo matching setup;
  3. 3.
    our method demonstrates state-of-the-art results on the Helvipad [8] dataset, a challenging real-world benchmark for omnidirectional stereo matching. In addition, DFI-OmniStereo shows generalization capabilities to other datasets and high sample efficiency.
@@ II RELATED WORK
A disparity map can be estimated from two images and then deterministically converted to depth assuming a calibrated stereo setup. Pre-deep learning approaches [12, 13] follow four steps: matching cost computation, cost aggregation, disparity computation, and refinement. Early deep learning methods [14] use Convolutional Neural Networks (CNNs) for matching cost calculation or end-to-end disparity estimation [15]. More recent 3D networks [16, 17, 18, 19] introduce 4D cost volumes via feature concatenation and employ a 3D encoder-decoder design for context aggregation. A large portion of recent work in deep stereo matching follows two paradigms: Vision Transformer [20] (ViT)-based methods and iterative optimization-based methods [21]. GMStereo [22] and CroCo-Stereo [23] propose a unified transformer-based architecture for optical flow and stereo. RAFT-Stereo [24] presents an iterative multi-resolution approach that refines disparity maps using a multi-level 3D correlation volume and a learned context encoding via convolutional GRUs [25]. Following RAFT-Stereo, many subsequent architectures [26, 27, 28, 29] propose an iterative refinement of the disparity. In particular, IGEV-Stereo [30] constructs a combined 3D cost volume that fuses local cues from a multi-level 3D correlation volume with global context from a regularized 4D cost volume to iteratively refine disparity estimates. ViTAStereo [31] utilizes features from the vision foundation model DINOv2 [32] in an iterative optimization-based method. Our approach extends this family by incorporating a large-scale pre-trained depth foundation model as feature extractor to enhance generalization. Unlike ViTAStereo, which uses only a foundation model-based encoder, we also fine-tune the decoder of a foundation model to (omnidirectional) stereo matching and consider the strong correlation between relative depth and disparity in our feature extractor selection.
@@
Three-dimensional scene geometry is inferred from data captured over a 360° field of view in omnidirectional depth estimation. Several works [33, 34, 35] approach this task by applying an equirectangular projection to map the spherical field of view of omnidirectional cameras onto a plane. However, this results in distortions at the top and bottom of the image [36]. To address these distortions, some works adapt CNNs to spherical images [37, 38, 33], while others develop distortion-aware ViTs [39, 40]. A second stream of works employs alternative projection methods [41, 42, 43], such as cubemaps [41]. Most omnidirectional stereo matching methods [6, 7, 9, 10, 34, 44, 45] rely on four fisheye cameras capturing images from different directions with overlapping fields of view, enabling processing via stereo matching. However, all these methods validate their results exclusively on synthetic datasets. A simpler and more cost-effective setup uses just two omnidirectional cameras, a top and a bottom one. However, until recently, research has been limited by the lack of large-scale datasets [8]. To the best of our knowledge, the only existing architectures for omnidirectional stereo matching with this configuration are 360SD-Net [46] and 360-IGEV-Stereo [8]. 360SD-Net, built on PSMNet [17], addresses distortions by encoding a polar angle image and concatenating it with the image features. A learnable vertical shifting filter is used to adjust for varying pixel step sizes in the cost volume construction. 360-IGEV-Stereo, built on IGEV-Stereo, integrates an encoded polar angle map into its feature and context networks. Additionally, it applies circular padding [47] before inference to exploit the circular boundary conditions and constructs cost volumes via vertical instead of horizontal warping to suit the camera setup. However, this previous approach yields inaccurate results near object boundaries under diverse lighting conditions (cf. Figure 3) due to the limited robustness of its feature network.
@@
Figure 2: Overview of DFI-OmniStereo. A shared depth foundation model (purple) is utilized to extract representations from a top and bottom image. Subsequently, an omnidirectional stereo matching head (pink) predicts disparity, utilizing the image features as follows: The intermediate representations and relative depth maps of both images are adapted to be processed as multi-scale feature maps by the iterative matching head. This head predicts a disparity map using vertical warping for cost volume construction. The training consists of two stages. In training stage A (yellow), we adapt the stereo matching head to the omnidirectional data and the foundation model features (foundation model frozen) using a conventional stereo matching loss $\mathcal{L}_{A}$. In stage B (orange), we fine-tune the foundation model decoder and the stereo matching head, utilizing a scale-invariant logarithmic loss $\mathcal{L}_{B}$. Frozen and trainable modules are denoted with a snowflake and fire symbol, respectively.
@@ IV-A Dataset and Evaluation Metrics
We train and evaluate our approach on Helvipad [8], the only real-world omnidirectional stereo depth estimation dataset with a top-bottom camera setup. The dataset consists of 27K training, 3K validation, and 10K test image pairs, along with ground-truth depth and disparity maps. Each split features a mix of indoor scenes, outdoor daytime, and outdoor nighttime scenes. Following [8], the task is defined as estimating disparity $d=\theta_{b}-\theta_{t}$ from a top-bottom stereo image pair, where $\theta_{t}$ and $\theta_{b}$ are the polar angles of the top and bottom cameras’ spherical camera model. According to [8] the disparity $d$ can be converted to depth $r_{b}$ using
@@
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@ IV-B Implementation Details
We implement our model in PyTorch [57], building upon the codebases of Depth Anything [^1] https://github.com/LiheYoung/Depth-Anything and IGEV-Stereo. [^2] https://github.com/gangweiX/IGEV To stabilize training, the disparity $d_{\text{deg}}$ in degrees is clamped at each iteration within the range $[d_{\text{deg, min}},d_{\text{deg, max}}]$ according to the dataset statistics. We convert the disparity, expressed in degrees, into pixels to properly warp the top image for constructing the cost volume, setting accordingly the maximum disparity for the cost volume to 128. The model is trained at full resolution 512$\times$1920, using standard photometric data augmentations.
For initialization, the unmodified components of the stereo matching head are initialized with IGEV-Stereo weights obtained from Scene Flow pre-training [15], while the modified components are initialized randomly. The image encoder ViT and the DPT decoder are initialized using the Depth-Anything-V2-Base checkpoint. [^3] https://github.com/DepthAnything/Depth-Anything-V2
During the feature adaptation stage, we use a batch size of $2$ for $20$ epochs with a learning rate of $2e^{-4}$. In the subsequent fine-tuning stage, we use a batch size of $1$ for $12$ epochs with a reduced learning rate of $2e^{-5}$. Additionally, we reduce the learning rate for the foundation model weights by a factor of $50$. In the losses, we set $\gamma=0.9$ and $\lambda=0.15$. Following the recommendations in [8], we use the depth-completed ground truth for training and evaluate all metrics except for the LRCE with the original sparse ground truth. Furthermore, we apply circular padding by $64$ pixels during inference. All experiments are conducted on a single NVIDIA A100 GPU.
@@ IV-C Comparison to State of the Art
Table I reports the comparative results on the Helvipad test split. DFI-OmniStereo outperforms the considered methods across nearly all evaluation metrics. Notably, our method achieves the lowest disparity MAE (0.158°), which is significantly lower than the next best result, 360-IGEV-Stereo (0.188°). The left-right consistency metric LRCE of our method is slightly worse than the best-performing baseline (360-IGEV-Stereo). We hypothesize that this outcome may stem from the inherent design of the ViT encoder and its positional encodings, which may not preserve local continuity as effectively as a CNN-based backbone does in capturing the transitions at the image boundaries after circular padding. Overall, the quantitative results of DFI-OmniStereo demonstrate that foundation model features can be leveraged to reduce errors and improve disparity and depth accuracy in omnidirectional stereo matching on diverse real-world data.
@@
TABLE II: DFI-OmniStereo component training analysis by selectively training different architectural components during the two training stages. All remaining components are frozen. We use the following abbreviations: FE (feature encoder), FD (feature decoder), and OS (omnidirectional stereo matching). indicates an adjusted learning rate due to unstable training.
@@ IV-D Analyzing DFI-OmniStereo
We analyze the impact of applying different loss terms in the two training stages. As shown in Table III, choosing the L1-based loss leads to the best results during the first training stage. Building on the better configuration, the SILog loss proves to be superior in the second training stage to adapt the relative depth representations to the omnidirectional imagery. These findings show the importance of starting from an L1-based loss before transitioning to a scale-invariant loss during training.
TABLE III: DFI-OmniStereo loss analysis. We explore the impact of using different loss terms across the two training stages. Stage B builds upon the best setup (L1-based) of the previous stage.
@@
Comparison to the depth foundation model
In Table IV, we compare DFI-OmniStereo to the monocular depth foundation model Depth Anything used within our architecture. Out of the box, this foundation model generalizes very poorly to omnidirectional images. [^4] We use Depth Anything V1 here, since checkpoints for the ZoeDepth [58] head fine-tuned for metric depth prediction are not available for V2. However, Yang et al. [56] show that this difference should not significantly impact accuracy. As an additional comparison, we extensively fine-tune Depth Anything V2 alongside the ZoeDepth [58] metric depth prediction head on the Helvipad dataset. Fine-tuning the model likely compromises its generalization properties but provides an “upper bound” of monocular depth estimation in this context. Notably, DFI-OmniStereo is consistently better across all metrics on the Helvipad dataset, highlighting its ability to leverage the stereo setup. The large disparity MAE difference suggests that our model excels at handling low depth values.
TABLE IV: Comparing DFI-OmniStereo to Depth Anything pre-trained and fine-tuned on metric depth prediction. indicates the use of the large-scale pre-trained checkpoint that includes the metric depth prediction head. refers to training using a ZoeDepth head [58] for metric depth estimation following the official Depth Anything code [11].
@@ IV-E Sample-efficient Learning
Collecting labeled real-world data is expensive. Having methods that can learn from a small amount of samples is essential for real-world applications. Figure 4 shows that DFI-OmniStereo already achieves a lower disparity MAE than 360-IGEV-Stereo (with 100% training data) when only $5\%$ of the training data, randomly sampled, are available to our model. This highlights how leveraging large-scale pre-training from a foundation model significantly reduces the need for task-specific data.
(a)    Disparity MAE
(b)    Disparity MARE
@@ IV-F Qualitative Generalization Analysis
Finally, we aim to assess the generalization capabilities of our method. 360SD-Net [46] is the only work beyond Helvipad providing real-world data with a similar top-bottom camera setup. The authors provide three scenes without ground-truth annotations. We qualitatively analyze the cross-dataset generalization of DFI-OmniStereo comparing to 360-IGEV-Stereo in Figure 5. Fine details, such as the chairs and table on the left of the hall scene, are only recognizable in DFI-OmniStereo’s disparity map. Our method predicts the depth boundaries of objects more accurately. For example, the brown chair and the blue desk chair in the room scene are only distinguishable in DFI-OmniStereo’s prediction. Homogeneous surfaces, such as the wall behind the left stairs in the stairs scene, are better visible in DFI-OmniStereo’s disparity prediction.
These results demonstrate promising generalization capabilities of our method DFI-OmniStereo when transferring to new cameras and stereo baselines.
@@
Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative comparison of generalization to real-world images from [46]. We visualize the bottom image and the disparity maps (°) output by 360-IGEV-Stereo and DFI-OmniStereo (from top to bottom) using the hall, room, and stairs scene (from left to right).
V CONCLUSION
In this work, we introduce DFI-OmniStereo, an omnidirectional stereo matching approach that integrates a large-scale pre-trained monocular depth foundation model into an iterative optimization-based stereo matching framework. Thanks to a two-stage training strategy, we ensure feature adaptation to omnidirectional stereo matching while preserving the generalization capabilities acquired by the foundation model during its pre-training. Extensive experiments on the Helvipad dataset demonstrate that DFI-OmniStereo outperforms the previous state of the art by a large margin in all depth and disparity metrics. Additionally, our model exhibits good generalization capabilities to unseen real-world images and is training sample efficient, highlighting its potential for real-world robotics applications. Future work could extend our approach of incorporating large-scale pre-trained depth foundation models to other highly specialized stereo settings like aerial and underwater imaging, where challenges further include varying illumination and water refraction.

### v2

@@ Abstract
Omnidirectional depth perception is essential for mobile robotics applications that require scene understanding across a full 360° field of view. Camera-based setups offer a cost-effective option by using stereo depth estimation to generate dense, high-resolution depth maps without relying on expensive active sensing. However, existing omnidirectional stereo matching approaches achieve only limited depth accuracy across diverse environments, depth ranges, and lighting conditions, due to the scarcity of real-world data. We present DFI-OmniStereo, a novel omnidirectional stereo matching method that leverages a large-scale pre-trained foundation model for relative monocular depth estimation within an iterative optimization-based stereo matching architecture. We introduce a dedicated two-stage training strategy to utilize the relative monocular depth features for our omnidirectional stereo matching before scale-invariant fine-tuning. DFI-OmniStereo achieves state-of-the-art results on the real-world Helvipad dataset, reducing disparity MAE by approximately 16% compared to the previous best omnidirectional stereo method.
@@ I INTRODUCTION
Figure 1: Overview of our proposed omnidirectional stereo matching approach DFI-OmniStereo. Given a pair of equirectangular images captured by two vertically stacked omnidirectional cameras, our method integrates a large-scale pre-trained monocular relative depth foundation model into an iterative stereo matching approach. DFI-OmniStereo improves disparity and depth estimation accuracy, significantly outperforming the previous state-of-the-art method. We visualize predicted disparity on a log-scale (red indicates high disparity and low depth; vice versa for blue).
@@
Specifically, our contributions are as follows: (1) We leverage a large-scale pre-trained monocular depth foundation model as a feature extractor integrated into an iterative optimization-based stereo matching architecture. (2) We design a two-stage training strategy to adapt monocular foundation model features to the omnidirectional stereo matching setup. We partially fine-tune the foundation model, and employ a scale-invariant error in log space (SILog loss) for stereo matching. (3) Our method demonstrates state-of-the-art results on the Helvipad [8] dataset, a challenging real-world benchmark for omnidirectional stereo matching. In addition, DFI-OmniStereo shows generalization capabilities to other datasets and high sample efficiency.
@@ II RELATED WORK
A disparity map can be estimated from two images and then deterministically converted to depth assuming a calibrated stereo setup. Early deep learning methods [12] use Convolutional Neural Networks (CNNs) for matching cost calculation or end-to-end disparity estimation [13]. More recent 3D networks [14, 15, 16, 17] introduce 4D cost volumes via feature concatenation and employ a 3D encoder-decoder design for context aggregation. Many recent works in deep stereo matching follow two paradigms: Vision Transformer [18] (ViT)-based methods and iterative optimization-based methods [19]. GMStereo [20] and CroCo-Stereo [21] propose a unified transformer-based architecture for optical flow and stereo. RAFT-Stereo [22] presents an iterative multi-resolution approach that refines disparity maps using a multi-level 3D correlation volume and a learned context encoding via convolutional GRUs [23]. Following RAFT-Stereo, many subsequent architectures [24, 25, 26, 27] propose an iterative refinement of the disparity. In particular, IGEV-Stereo [28] constructs a 3D cost volume that fuses local cues from a multi-level 3D correlation volume with global context from a regularized 4D cost volume to iteratively refine disparity estimates. Vankadar et al. [29] and ViTAStereo [30] use features from the vision foundation models DINO [31] and DINOv2 [32] for stereo matching. Our approach extends this family by incorporating a large-scale pre-trained depth foundation model as feature extractor to consider the strong correlation between relative depth and disparity. Unlike concurrent work,111 Note that, independently and concurrently with our work, Wen et al. [33] and Cheng et al. [34] adopt a conceptually similar methodology to DFI-OmniStereo, applying it to conventional stereo matching. We also fine-tune the decoder of a foundation model to (omnidirectional) stereo matching.
@@
Three-dimensional scene geometry is inferred from data captured over a 360° field of view in omnidirectional depth estimation. Several works [35, 36, 37] approach this task by applying an equirectangular projection to map the spherical field of view of omnidirectional cameras onto a plane. However, this results in distortions at the top and bottom of the image [38]. To address these distortions, some works adapt CNNs to spherical images [39, 40, 35], while others develop distortion-aware ViTs [41, 42, 43]. A second stream of works employs alternative projection methods [44, 45, 46], such as cubemaps [44]. Most omnidirectional stereo matching methods [6, 7, 9, 10, 36, 47, 48] rely on four fisheye cameras capturing images from different directions with overlapping fields of view, enabling stereo matching. However, all these methods validate their results exclusively on synthetic datasets. A simpler and more cost-effective setup uses just two omnidirectional cameras, a top and a bottom one. However, until recently, research has been limited by the lack of large-scale datasets [8]. To the best of our knowledge, the only existing architectures for omnidirectional stereo matching with this configuration are 360SD-Net [49] and 360-IGEV-Stereo [8]. 360SD-Net, built on PSMNet [15], addresses distortions by encoding a polar angle image and concatenating it with the image features. A learnable vertical shifting filter is used to adjust for varying pixel step sizes in the cost volume construction. 360-IGEV-Stereo, built on IGEV-Stereo, integrates an encoded polar angle map into its feature and context networks. Additionally, it applies circular padding [50] before inference to exploit the circular boundary conditions and constructs cost volumes via vertical instead of horizontal warping to suit the camera setup. However, this previous approach yields inaccurate results near object boundaries under diverse lighting conditions (cf. Figure 3) due to the limited robustness of its feature network.
@@
Figure 2: Overview of DFI-OmniStereo. A shared depth foundation model (purple) is utilized to extract representations from a top and bottom image. Subsequently, an omnidirectional stereo matching head (pink) predicts disparity, utilizing the image features as follows: The intermediate representations and relative depth maps of both images are adapted to be processed as multi-scale feature maps by the iterative matching head. This head predicts a disparity map using vertical warping for cost volume construction. The training consists of two stages. In training stage A (blue), we adapt the stereo matching head to the omnidirectional data and the foundation model features (foundation model frozen) using a conventional stereo matching loss $\mathcal{L}_{A}$. In stage B (orange), we fine-tune the foundation model decoder and the stereo matching head, utilizing a scale-invariant logarithmic loss $\mathcal{L}_{B}$. Frozen and trainable modules are denoted with a snowflake and fire symbol, respectively.
@@ IV-A Dataset and Evaluation Metrics
We train and evaluate our approach on Helvipad [8], the only real-world omnidirectional stereo depth estimation dataset with a top-bottom camera setup. While there exist synthetic datasets, Stereo-MP3D [49, 60] and Stereo-SF3D [49, 61], which share our configuration, their lack of photorealism makes them unsuitable for training and evaluation. Wang et al. [49] additionally provide three real-world images without ground truth, which we include in a qualitative analysis in Section IV-F. The Helvipad dataset consists of 27K training, 3K validation, and 10K test image pairs, along with ground-truth depth and disparity maps. Each split features a mix of indoor scenes, outdoor daytime, and outdoor nighttime scenes. Following [8], the task is defined as estimating disparity $d=\theta_{b}-\theta_{t}$ from a top-bottom stereo image pair, where $\theta_{t}$ and $\theta_{b}$ are the polar angles of the top and bottom cameras’ spherical camera model. According to [8] the disparity $d$ can be converted to depth $r_{b}$ using
@@
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
@@
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
@@
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
$^{-1}1001$$^{0}$$10^{1}$ $^{-1}1001$$^{0}$$10^{1}$ $^{-1}1001$$^{0}$$10^{1}$
@@ IV-B Implementation Details
[…]
@@ IV-C Comparison to State of the Art
Table I reports the comparative results on the Helvipad test split. DFI-OmniStereo outperforms the considered methods across nearly all evaluation metrics. Notably, our method achieves the lowest disparity MAE (0.158°), which is significantly lower than the next best result, 360-IGEV-Stereo (0.188°). The left-right consistency metric LRCE of our method is marginally worse than the best-performing baseline (360-IGEV-Stereo). Overall, the quantitative results of DFI-OmniStereo demonstrate that foundation model features can be leveraged to reduce errors and improve disparity and depth accuracy in omnidirectional stereo matching on diverse real-world data.
@@
TABLE II: DFI-OmniStereo component analysis by selectively training different architectural components during the two training stages. All remaining components are frozen. We use the following abbreviations: FE (feature encoder), FD (feature decoder), and OS (omnidirectional stereo matching). indicates an adjusted learning rate due to unstable training.
@@ IV-D Analyzing DFI-OmniStereo
We analyze the impact of applying different loss terms in the two training stages. As shown in Table III, choosing the L1-based loss leads to the best results during the first training stage. Building on the better configuration, the SILog loss proves to be superior in the second training stage to adapt the relative depth representations to the omnidirectional imagery. These findings indicate the importance of starting with an L1-based loss before transitioning to a scale-invariant loss.
TABLE III: DFI-OmniStereo loss analysis. We explore the impact of using different loss terms across the two training stages. Stage B uses the best setup (L1-based) of stage A.
@@
Comparison to monocular depth estimation models
In Table IV, we compare DFI-OmniStereo to the monocular depth estimation methods EGformer [42], Depth Anything [11], and Depth Anything V2 [59]. EGformer [42] is a monocular relative depth estimation approach specialized for omnidirectional imagery. We adapt EGformer for metric depth estimation by replacing the final activation, removing the scale-and-shift alignment of the loss, and adjusting the spherical grid to match the Helvipad images. Despite being designed for equirectangular images, EGformer performs significantly worse than DFI-OmniStereo across all metrics. Out of the box, the Depth Anything model generalizes very poorly to omnidirectional images. [^2] We use Depth Anything V1 here, since checkpoints for the ZoeDepth [63] head fine-tuned for metric depth prediction are not available for V2. However, Yang et al. [59] show that this difference should not significantly impact accuracy. In addition, we extensively fine-tune Depth Anything V2 alongside the ZoeDepth [63] metric depth prediction head on the Helvipad dataset. DFI-OmniStereo is consistently better across all metrics on the Helvipad dataset, highlighting the benefit of leveraging stereo cues in combination with the depth foundation model features. We further analyze where DFI-OmniStereo improves over Depth Anything V2 by comparing disparity MAE across three depth intervals in Table V. Each interval comprises around one third of the available ground-truth values. Notably, our method performs particularly well at medium depth ranges ($4\text{\,}\mathrm{m}$$9\text{\,}\mathrm{m}$) with a disparity MAE reduction of $8.7\%$.
TABLE IV: Comparison of DFI-OmniStereo to monocular depth estimation models. All models are adapted to metric depth estimation. $\ast$ refers to our adapted version for metric depth estimation. $\dagger$ indicates the use of the large-scale pre-trained checkpoint that includes the metric depth prediction head. $\ddagger$ refers to training using a ZoeDepth head [63] for metric depth estimation following [11].
@@
TABLE V: Comparing DFI-OmniStereo to Depth Anything for different depth ranges. We explore the disparity MAE (in °) for three depth ranges (in  $\text{\,}\mathrm{m}$).
[…]
Depth Anything V2 [59] $0.182$ $0.149$ $0.148$
DFI-OmniStereo $0.181$ $0.136$ $0.150$
@@ IV-E Sample-efficient Learning
Collecting labeled real-world data is expensive. Having methods that can learn from a small amount of samples is essential for real-world applications. Figure 4 shows that DFI-OmniStereo already achieves a lower disparity MAE than 360-IGEV-Stereo (with $100\%$ training data) when only $5\%$ of the training data, randomly sampled, are available to our model. This highlights how leveraging large-scale pre-training from a foundation model significantly reduces the need for task-specific data.
1251020501000.160.160.180.180.20.20.220.22Sampling ratio (%)Disp-MAE ()Disparity MAE360-IGEV-Stereo (100%)DFI-OmniStereo (Ours) 1251020501000.120.120.140.140.160.16Sampling ratio (%)Disp-MAREDisparity MARE360-IGEV-Stereo (100%)DFI-OmniStereo (Ours)
@@ IV-F Qualitative Generalization Analysis
Finally, we aim to assess the generalization capabilities of our method. 360SD-Net [49] is the only work beyond Helvipad providing real-world data with a similar top-bottom camera setup. The authors provide three scenes without ground-truth annotations. We qualitatively analyze the cross-dataset generalization of DFI-OmniStereo comparing to 360-IGEV-Stereo in Figure 5. Fine details, such as the chairs and table on the left of the hall scene, are only recognizable in DFI-OmniStereo’s disparity map. Our method predicts the depth boundaries of objects more accurately. For example, the brown chair and the blue desk chair in the room scene are only distinguishable in DFI-OmniStereo’s prediction. Homogeneous surfaces, such as the wall behind the left stairs in the stairs scene, are better visible in DFI-OmniStereo’s disparity prediction. These results demonstrate promising generalization capabilities of our method DFI-OmniStereo when transferring to new cameras and stereo baselines.
@@
$^{-1}1001$$^{0}$ $^{-1}1001$$^{0}$$10^{1}$ $^{-1}1001$$^{0}$
Figure 5: Qualitative comparison of generalization to real-world images from [49]. We visualize the bottom image and the disparity prediction (°) of 360-IGEV-Stereo and DFI-OmniStereo (from top to bottom) using the hall, room, and stairs scene (from left to right).
V LIMITATIONS AND FUTURE WORK
While DFI-OmniStereo achieves state-of-the-art metric results, we did not focus on efficiency and real-time capabilities. Future work could address this through model compression techniques, knowledge distillation into a smaller task-specific foundation model component, or by replacing the iterative stereo matching. We rely on the only real-world omnidirectional stereo depth dataset with a top-bottom camera setup, so we test the large pre-trained depth model in a single specialized setting. Future work should evaluate its generalization to other data-scarce stereo scenarios, such as aerial, or underwater imaging.
VI CONCLUSION
We introduced DFI-OmniStereo, an omnidirectional stereo matching approach that integrates a large-scale pre-trained monocular depth foundation model into an iterative optimization-based stereo matching framework. Thanks to a two-stage training strategy, we ensure feature adaptation to omnidirectional stereo matching while preserving the generalization capabilities acquired by the foundation model during its pre-training. Extensive experiments on the Helvipad dataset demonstrate that DFI-OmniStereo outperforms the previous state of the art by a large margin across multiple depth and disparity metrics. Additionally, our model exhibits good generalization capabilities to unseen real-world images and is training sample efficient, highlighting its potential for real-world robotics applications.